OpenStudy (anonymous):

identify the vertical asymptotes of f(x)=2/x^2+3x-10

OpenStudy (anonymous):

@hartnn Is that right? ^

OpenStudy (anonymous):

i think its -5

hartnn (hartnn):

yes, but where does that other f(x) come from ? and why that oblique asymptote ? are you copying answers from other site ?

hartnn (hartnn):

x=2 and x=-5 are correct.

OpenStudy (anonymous):

@hartnn Oh, sorry I was answering two problems. My notepad is a little messy sorry. Let me fix that.

OpenStudy (anonymous):

f(x) = 2/(x² + 3x - 10) = 2 / [(x + 5)(x - 2)] Vertical asymptotes are x=2 and x=-5.

OpenStudy (anonymous):

thnk you

hartnn (hartnn):

vertical asymptotes are when denominator goes to 0

OpenStudy (anonymous):

@helpineedit Your welcome :)